OpenStudy (anonymous):

A lion's heart beats 12 times in 16 seconds. How many heartbeats will it have in 60 seconds? A) 3.2 heartbeats B) 36 heartbeats C) 45 heartbeats D) 60 heartbeats

OpenStudy (anonymous):

@ChiefArnav

OpenStudy (anonymous):

fan and medal for best answer

OpenStudy (nurali):

12/16= 3/4 x/60=3/4 Cross multiply. 4x=180 Divide by 4 x=45

OpenStudy (anonymous):

beats time 12 16 x 60

OpenStudy (anonymous):

thank you guys sooooooooo much

OpenStudy (anonymous):

you mutiply 12 by 4 and you get 48 so 48 beats for 60 seconds

OpenStudy (anonymous):

Oops nvm its 45 16 is by 3 not 4 lol